# Experiment Study on Reference-Path-Aided System Calibration for mmWave Bistatic ISAC Systems

### v1

@@ I Introduction
There are two basic modes to integrate sensing function with communication networks: monotonic ISAC and bistatic ISAC. In monostatic ISAC, the transmitter and sensing receiver are co-located or integrated in the same transceiver. Therefore, the full-duplex radio serves as the key enabler, with self-interference cancellation being the primary challenge [3]. In bistatic ISAC, the transmitter and sensing receiver are separate devices and usually deployed at different positions, e.g., they can be two different base stations (BS). Half-duplex radios can be utilized to support bisatic sensing. Given that existing communication network infrastructures are based entirely on half-duplex radios, bistatic sensing emerges as a promising approach for broad deployment in future ISAC networks, allowing the sensing receiver to fully leverage the communication receiver hardware for wireless sensing. For example, the wireless sensing function can be achieved by processing the received OFDM symbols or channel state information (CSI) [4]. However, since the transmitter and receiver are separate devices, the synchronization between the two devices becomes a key challenge for wireless sensing [5].
@@ II-A System Model
In this paper, a bistatic ISAC system is considered, where the transmitter (Tx) and sensing receiver (Rx) localizations are fixed. For example, the Tx and sensing Rx can be two BSs in B5G/6G networks. In practice, the direct link, i.e., the LoS path between the Tx and Rx, may experience intermittent blockage due to moving obstacles, as shown in Fig. 1. Therefore, there may either be a LoS path between the transmitter and sensing receiver, or only NLoS paths between them. For the LoS scenario, the LoS path can be utilized as a reference path for system calibration. To support the NLoS scenario, a strong NLoS path is required for system calibration, which can be found in the environment or manually deployed. For instance, a RIS can be deployed as a reference reflector to form a strong NLoS path. The orientations and locations of the Tx, Rx, and reference reflector are prior knowledge. Therefore, in the LoS scenario, the angle of arrival (AoA) of the LoS path and the distance between the Tx and Rx are provided, whereas in the NLoS scenario, the AoA of the reference path and the combined distance along the Tx–reference reflector–Rx path are given.

### v2

# Experimental Study on Reference-Path-Aided System Calibration for mmWave Bistatic ISAC Systems

@@ I Introduction
There are two basic modes to integrate sensing function with communication networks: monotonic ISAC and bistatic ISAC. In monostatic ISAC, the transmitter and sensing receiver are co-located or integrated in the same transceiver. Therefore, the full-duplex radio serves as the key enabler, with self-interference cancellation being the primary challenge [3]. In bistatic ISAC, the transmitter and sensing receiver are separate devices and usually deployed at different positions, e.g., they can be two different base stations (BS). Half-duplex radios can be utilized to support bisatic sensing. Given that existing communication network infrastructures are based entirely on half-duplex radios, bistatic sensing emerges as a promising approach for broad deployment in future ISAC networks. Since sensing and communication share the same physical channel in bistatic ISAC systems[4], the communication receiver’s hardware can be fully reused for wireless sensing. For example, the wireless sensing can be achieved by processing the received OFDM symbols or channel state information (CSI) [5]. However, since the transmitter and receiver are separate devices, the synchronization between the two devices becomes a key challenge for wireless sensing [6].
@@ II-A System Model
In this paper, a bistatic ISAC system is considered, where the transmitter (Tx) and sensing receiver (Rx) localizations are fixed. For example, the Tx and sensing Rx can be two BSs in B5G/6G networks. In practice, the direct link, i.e., the LoS path between the Tx and Rx, may experience intermittent blockage due to moving obstacles, as shown in Fig. 1. Therefore, there may either be a LoS path between the transmitter and sensing receiver, or only NLoS paths between them. For the LoS scenario, the LoS path can be utilized as a reference path for system calibration. To support the NLoS scenario, a strong NLoS path is required for system calibration, which can be found in the environment or manually deployed. For instance, a RIS can be deployed as a reference reflector to form a strong NLoS path. The orientations and locations of the Tx, Rx, and reference reflector are prior knowledge. Therefore, in the LoS scenario, the angle of arrival (AoA) of the LoS path and the distance between the Tx and Rx are provided, whereas in the NLoS scenario, the AoA of the reference path and the combined distance along the Tx-reference reflector-Rx path are given.